Average price of a CD reduces from $21 to $15. Company expected the quantity will increase 30%.

a. calculate the price elasticity of demand over the price range. 


1
Expert's answer
2021-01-12T13:33:54-0500

price elasticity of demand "= \%" change in quantity "\/" "\%" change in price

"\%" change in price "=\frac{15-21}{21}\times 100\%=-28.5714\%"

price elasticity of demand"=30\% /-28.5714\%=-1.05=|-1.05|=1.05"